OpenStudy (anonymous):

Danny’s 7 friends contributed an equal amount of money to buy him a birthday present. The present cost $84. How much money did each friend contribute?

jimthompson5910 (jim_thompson5910):

Each friend contributed 84/7 = 12 dollars

jimthompson5910 (jim_thompson5910):

You can check your answer by multiplying the cost per person (12) by the number of people (7) to get 12*7 = 84

OpenStudy (anonymous):

Ok thank you, Cause used addition.. Which was wayyyy longer

jimthompson5910 (jim_thompson5910):

You can do that too, you can say 12+12+12+12 + 12+12+12 = 84 but yes, it's much longer. If you can remember that multiplication is a shortcut for repeated addition, then you'll save a lot of time.

jimthompson5910 (jim_thompson5910):

Oh and division is the opposite of multiplication, which helps out a lot as well.